OpenStudy (anonymous):

3/10 divided by 7/15

OpenStudy (anonymous):

0.00285714285

OpenStudy (anonymous):

flip the 2nd faction around and multiply

OpenStudy (zarkon):

\[\frac{3}{10}\div\frac{7}{15}=\frac{\frac{3}{10}}{\frac{7}{15}}=\frac{\frac{3}{10}\cdot\frac{15}{7}}{\frac{7}{15}\cdot\frac{15}{7}}=\frac{\frac{3}{10}\cdot\frac{15}{7}}{1}=\frac{3\cdot15}{10\cdot7}=\cdots\]

OpenStudy (anonymous):